Plzz..ans. this question..[calculate the pH of a mixture containing 50 mililitre of 1(N) HCl and 30 mililitre of 1(N) NaOH].((Ans. ....0.6)).

Plzz..ans.  this question..[calculate the pH of a mixture containing 50 mililitre  of 1(N) HCl and 30 mililitre of 1(N) NaOH].((Ans. ....0.6)).

Grade:11

1 Answers

Arun
5 years ago

Mole = concentration x volume

mole of NaOH = 30ml x 1 = 30.0

mole of HCl = 50ml x 1= 50.0

30 mole of NaOH exactly neutralized by 30 mole of HCl.

NaOH + HCl = NaCl + H2O

that is 50 -30 = 20 HCl mole

Con of H+ = no of mole of H+ / total volume of solution (50+ 30 ml = 80 ml)

= 20/80 = 0.25

pH = - log ( H+)

= - log (1/4) = 0.6021

s0 the pH of solution will be 0.6021